# On the geometric phase and its role in the design of elastic topological materials

### v1

@@ 4.2 Connections and parallel transport
Figure 5 illustrates parallel transport on the fiber bundle of Fig. 3 using two (contrived) connections. Let $C$ be a curve and $L$ be a loop on the base manifold. Figure 5a depicts a connection where parallel transport of an element $u_{1}$ of the fiber bundle along $C$ does not change the fiber coordinate $f_{1}$. Such a connection is said to be flat. It follows that parallel transport of an element $u_{2}$ of the fiber bundle along the loop $L$ does not change it. Figure 5b depicts a connection where parallel transport of an element $u_{1}$ of the fiber bundle along $C$ changes the fiber coordinate from $f_{1}$ to $f_{1}^{\prime}$. Such a connection is said to be curved. In this case, after parallel transport of an element $u_{2}$ of the fiber bundle along the loop $L$, the fiber coordinate generally does not return to its initial value. The change in the fiber element $\gamma$ is known as a holonomy.
Figure 5: Parallel transport on the fiber bundle of Fig. 3 according to a connection. (a) Fiber bundle with a trivial connection. (b) Fiber bundle with a curved connection.
@@
Figure 6a illustrates parallel transport on a bundle constructed by attaching a perpendicular line $\mathbb{R}$ at each point on the plane $\mathbb{R}^{2}$. The bundle is known as the normal bundle over the plane, and it is denoted as $\mathcal{N}\mathbb{R}^{2}$. Its base manifold is $\mathbb{R}^{2}$, and the fiber is $\mathbb{R}$. As the base manifold is a Euclidean space, $\mathcal{N}\mathbb{R}^{2}$ is trivial. ($\mathcal{N}\mathbb{R}^{2}$ is, in fact, equal to $\mathbb{R}^{3}$ by construction.) A point $u$ of $\mathcal{N}\mathbb{R}^{2}$ describes a vector $v$ normal to the plane $\mathbb{R}^{2}$ that is attached at a point $\mathbf{R}$ of $\mathbb{R}^{2}$. Its fiber coordinate is the vector $v$ and the base coordinate is the point $\mathbf{R}$. Parallel transport on $\mathcal{N}\mathbb{R}^{2}$ corresponds to moving normal vectors along a curve on the plane without rotating them or changing their lengths. The fiber coordinate does not change along the path upon parallel transport (Fig. 6a), indicating that $\mathcal{N}\mathbb{R}^{2}$ is flat. The parallel transport of the fiber elements along any closed loop such as $\mathbf{R}_{1}\mathbf{R}_{2}\mathbf{R}_{3}\mathbf{R}_{4}$ does not result in a holonomy.
@@ 5.1 Effect on the overall phase
Consider a quasi-1D semi-infinite waveguide $W$ whose cross sectional properties are described by a set of $n$ parameters, $\mathbf{R}(z)$ (Fig. 7a). The parameters vary adiabatically in a cycle along the longitudinal direction. The space of the parameter values is generally equivalent to $\mathbb{R}^{n}$. Each parameter value $\mathbf{R}$ can define a uniform waveguide and its guided modes. Recall from Sec. 3 that a parameter value $\mathbf{R}_{0}$ resulting in a waveguide with fully degenerate guided modes is said to be a full degeneracy. The parameter space generally consists of nondegenerate values; only specific parameter values $P$ are fully degenerate. The nondegenerate parameter values defines the base manifold $\mathcal{M}$, that is, $\mathcal{M}=\mathbb{R}^{n}-P$ (Figs. 7b-d). A parameter space without full degeneracies results in a Euclidean base manifold and leads to a trivial fiber bundle (Fig. 7b). A parameter space with full degeneracies generally leads to a nontrivial fiber bundle (Fig. 7d).
@@
Figure 7: The geometric phase in an elastic waveguide is modeled using a fiber bundle. (a) A quasi-one-dimensional waveguide $W$ with cross sectional properties $\mathbf{R}$ varying in an adiabatic cycle along $z$. The waveguide is excited by a guided mode displacement amplitude $\mathbf{U(x,y,\mathbf{R}(0))}$ at $z=0$. The disturbance propagates to $z=z_{0}$, where the response acquires a geometric phase $\gamma$. (b) Base manifold $\mathcal{M}=\mathbb{R}^{2}$ and fiber $U(1)$ resulting in the fiber bundle $\mathcal{E}_{1}$. Waveguide $W$ corresponds to the loop $L$ on the base manifold, which is parameterized by the $z$ coordinate. The fiber $U(1)$ is represented by a line with identical endpoints. The color denotes the fiber coordinate from 00 to $2\pi$. The fiber is attached at each point, as shown at $\mathbf{R}(z_{0})$, to create the fiber bundle $\mathcal{E}_{1}$. Parallel transport of $u$ along $L$ models wave propagation in $W$. It results in a holonomy $\gamma$, which is a nontopological geometric phase. (c) Base manifold $\mathcal{M}=\mathbb{R}^{2}-\{\mathbf{0}\}$ and fiber $\{0,\pi\}$ resulting in a trivial fiber bundle $\mathcal{E}_{2}$. The fiber $\{0,\pi\}$, depicted by blue cubes, is attached at a point $\mathbf{R}(z_{0})$. (d) Base manifold $\mathcal{M}=\mathbb{R}^{2}-\{\mathbf{0}\}$ and fiber $\{0,\pi\}$ resulting in a nontrivial fiber bundle $\mathcal{E}_{3}$. Two loops $L_{1}$ and $L_{2}$ are marked. Parallel transport of an element $u$ along $L_{1}$ ($L_{2}$) results in a topological geometric phase (no geometric phase). (e) The edge of a Möbius strip, obtained by restricting $\mathcal{E}_{3}$ to loop $L_{1}$. Parallel transport of $u$ is highlighted. (f) The edges of a cylinder, obtained from restricting $\mathcal{E}_{3}$ to loop $L_{2}$. Parallel transport of $u$ is highlighted.
@@
The waveguide $W$ corresponds to a closed loop $L$ on the base manifold, which is parameterized by the longitudinal coordinate $z$ (Fig. 7b). Let the waveguide be harmonically excited with angular frequency $\omega$ triggering a flexural mode whose displacement amplitude is $\mathbf{U}(x,y;\mathbf{R}(0))$ at $z=0$. The steady state displacement field at a location $z_{0}$ is determined by the WKB approximation up to the geometric phase as (Sec. 3)
@@
The geometric phase $\gamma$ accounts for the phase ambiguity of the displacement amplitude (Sec. 2). That is, if $\mathbf{U}(x,y;\mathbf{R}(z_{0}))$ is a displacement amplitude function, then so is $\mathbf{U}(x,y;\mathbf{R}(z_{0}))e^{\mathrm{i}\gamma}$. In general, the geometric phase $\gamma$ may assume any value from 00 to $2\pi$, that is, any point on the complex unit circle $U(1)$. The geometric phase is described as the fiber coordinate on the fiber $U(1)$ attached to the point $\mathbf{R}$ of the base manifold. This is shown in Fig. 7b where the fiber is visualized as a line with the ends of the line identified as a single point and with the color denoting the fiber coordinate. The base manifold and fibers give rise to a fiber bundle (Fig. 7b). In systems with time reversal symmetry, such as conservative elastic waveguides, the guided modes are described by real-valued displacement amplitude functions. Thus, a computed displacement amplitude is ambiguous only in its sign, and the geometric phase is restricted to $\{0,\pi\}$. The geometric phase is described by the fiber $\{0,\pi\}$. Some possible fiber bundles are shown in Figs. 7c,d. In this manner, the response of a waveguide to a guided mode excitation has been modeled as a fiber bundle with base manifold as the space of nondegenerate parameters $\mathbf{R}$ and the fiber as $U(1)$ or $\{0,\pi\}$.
A connection on this fiber bundle is provided by the governing equations of motion and the adiabatic theorem. The adiabatic theorem limits the allowable magnitude of change in the displacement amplitude function over a small longitudinal distance $\Delta z$ to be of the order of $(\Delta z)^{2}$. That is,
@@
Equation (14) defines a connection called the adiabatic connection or the Berry connection [1, 14]. If the fiber is $\{0,\pi\}$, the adiabatic connection simply ensures that the fiber coordinate does not change abruptly (from 00 to $\pi$ or vice versa) as the base point moves in a neighborhood equivalent to the Euclidean space, implying that the connection is flat.
@@
Examples of such geometric phases in elastic waves can be found in Refs. [41, 37].
If the fiber bundle is nontrivial with fiber $\{0,\pi\}$, the geometric phase is topological. This situation is exemplified by the discussions on flexural waves propagating in 1D waveguides with adiabatically-varying triangular cross section (Sec. 3.1). The waveguide property varying along the longitudinal direction is a geometric perturbation $\mathbf{R}=(\Delta x,\Delta y)$ to the cross section. The perturbation parameters lead to a two-dimensional parameter space with a full degeneracy at $\mathbf{R}=(0,0)$ arising from the symmetry of the corresponding uniform waveguide, which has an equilateral triangular cross section. Thus, the base manifold is the punctured plane $\mathbb{R}^{2}-\{\mathbf{0}\}$, and the fiber is $\{0,\pi\}$. The base manifold and fiber gives rise one of the following fiber bundles: (i) a stack of two punctured planes $\mathcal{E}_{2}$ (formally $(\mathbb{R}^{2}-\{\mathbf{0}\})\times\{0,\pi\}$) shown in Fig. 7c, which is trivial, or (ii) a thickened Möbius strip-like bundle $\mathcal{E}_{3}$ (nontrivial) visualized in Fig. 7c (c.f. Sec. 4.1). Note that the self-intersection is an artifact of describing the fiber bundle in 3D space. As the simulations of Sec. 3.1 confirm the appearance of a $\pi$-valued geometric phase, the fiber bundle must be nontrivial (because a trivial fiber bundle with flat connection does not exhibit holonomy).
@@ 5.2 Effect on the polarization angle
For this case, the construction of the fiber bundle and identification of the connection is discussed using the example of polarized flexural waves propagating in a waveguide (Sec. 3.2).
Consider a waveguide $W$ generated by the closed loop $C_{2}$ on the sphere (Fig. 2g,h). Let the path coordinate along the centerline of the waveguide be $s$. Let the waveguide $W$ be excited by a flexural guided mode, assumed to be horizontally polarized. At any location $s$, the WKB approximation determines the displacement field up to the polarization vector $\hat{\mathbf{c}}$ (Sec. 3), which can be any unit vector in $\mathbb{R}^{2}$. It is modeled as the fiber coordinate on the fiber $\mathbb{R}^{2}$ attached to the point $\hat{\mathbf{t}}$ of the sphere. The resulting fiber bundle is equivalent to the tangent bundle of the sphere $TS^{2}$ (Fig. 3e).
@@ 5.3 General considerations
The previous two sections highlight a few general considerations for a waveguide $W$ with adiabatically varying parameters $\mathbf{R}$ along the propagation coordinate $z$ (or $s$): (i) the propagation of a harmonic excitation with displacement amplitude $\mathbf{U}(x,y,\mathbf{R}(0))$ is described as a parallel transport on a fiber bundle, (ii) the geometric phase is the holonomy arising from this parallel transport. The classification of holonomies according to the nature of the fiber bundle (trivial or nontrivial) and curvature of the connection (flat or curved) applies to geometric phases. In particular, a waveguide exhibiting a topological geometric phase has full degeneracies in the parameter space and is described by a nontrivial fiber bundle analogous to the edge of a Möbius strip. A waveguide exhibiting a nontopological geometric phase is described by a fiber bundle with a curved connection. This classification is verified in the examples in Sec. 3.
@@ 6 Topological classification of elastic waveguides
The discussions on geometric phase have illustrated the role of differential geometry and topology in understanding the dynamics of an elastic waveguide. In this section, we show that the relation between geometric phase, topology, and dynamic response can be leveraged to design elastic waveguides.
We review the role of the geometric phase in the design of a class of elastic waveguides called elastic metamaterials. Broadly speaking, elastic metamaterials [54, 55] are periodic structures. Their smallest repeating structural element is called a unit cell. Crucially, metamaterials derive their mechanical properties from the design of the unit cell, in addition to the material properties. The unit cell geometry can be tailored to enable a wide range of unconventional static and dynamic responses, such as negative Poisson’s ratio [56], negative refractive index [57], nonreciprocity [58], and cloaking [59, 60]. However, it is challenging to rationally navigate the vast design space to achieve a specific material response. Topological concepts provide guidance by classifying different unit cell designs based on their dynamical response.
[…]
The topological classification of abstract spaces and physical classification of materials were unified in a remarkable chain of discoveries in condensed matter physics between 1970 and 2000 (see a review in Refs. [31, 33]). These discoveries modeled crystalline materials as fiber bundles and linked physical properties to topological invariants. As an example, electric polarization was understood as the holonomy of parallel transport [18]. These investigations led to a “periodic table” of topological materials’ [33]. According to this classification scheme, given a material and its quantum mechanical symmetries, a topological invariant can be computed. Materials with the same invariant form a material class. The physical consequence of this classification scheme is visible at an interface of two materials with different invariants (or, equivalently, different topological behavior). These interfaces admit localized excitations, called edge modes, relevant for a variety of technological advances in electronics and information processing [63].
@@
In the last decades, some of these concepts as well as the possibility for topological classification of classical (i.e. non-quantum) materials has been explored and extended to elastic waves, acoustics, and photonics [25]. In the context of elastic metamaterials, research focused on the identification and realization of material architectures capable of mimicking the dynamic behavior of their quantum mechanical counterparts. While it is important to stress that foundational difference between quantum and classical materials (whose behavior is ruled by quantum and classical mechanics, respectively), which ultimately allows only analogies between the two material systems, studies indicated that the topological classification could be adapted to elastic metamaterials.
@@
Metamaterials designed using such topological ideas led to the field of elastic topological metamaterials (ETMs) that were characterized by very distinctive and unique wave propagation properties. An elastic metamaterial can be labeled with a topological invariant, which indicates certain characteristics of its dynamic response. Most notably, interfaces between topologically distinct elastic materials support localized propagating waves that are immune to back scattering even in the presence of inhomogeneities and defects. The initial development of ETMs was rapid since most designs emulated the design characteristics of their quantum counterparts. Nevertheless, researchers have only begun exploring ETM designs and capabilities [64, 65]. To support the development of ETMs, design tools rooted in differential geometry and topology are critical. However, currently these ideas are spread out over different fields and making their role in the design process of ETMs not always clear. This section represents a first attempt to fill this gap by using the notions introduced in Sec. 4 and provide a topological classification of ETMs.
@@
Consider the adiabatically-varying triangular waveguides presented in Sec. 3.1. A waveguide $W$ can be visualized as a closed loop, say $L_{1}$, on the base manifold $\mathbb{R}^{2}-\{\mathbf{0}\}$ (Fig. 8a), as discussed earlier. $L_{1}$ is parameterized by the longitudinal coordinate $z$.
A waveguide can also be visualized in terms of the dynamical properties of its cross section. Each point $\mathbf{R}$ on the loop $L_{1}$ corresponds to a cross section of the waveguide $W$ at some coordinate $z$ along the longitudinal axis. This cross section, with parameters $\mathbf{R}(z)$, defines a uniform waveguide denoted $W_{\mathbf{R}(z)}$. The dynamical properties of the flexural guided modes of the uniform waveguide $W_{\mathbf{R}(z)}$ are described by the dispersion relations $\omega(k,\mathbf{R}(z))$. By allowing $z$ to vary, $\omega(k,\mathbf{R}(z))$ can capture the dispersion relations of every cross section of $W$. For a convenient visualization, we fix $k$ and plot $\omega$ as a function of $z$ in Fig. 8b-d. This plot is referred to as the ‘$\omega$-$z$’ plot here. The two curves in an $\omega$-$z$ plot generally do not intersect. The curves intersect if and only if the flexural guided modes for some $W_{\mathbf{R}(z)}$ are fully degenerate, which means that the loop corresponding to $W$ passes through the full degeneracy (the origin). This situation is illustrated by loop $L_{2}$ in Fig. 8a and the corresponding dispersion curve in Fig. 8c.

### v3

@@ 4.2 Connections and parallel transport
Figure 5 illustrates parallel transport on the fiber bundle of Fig. 3 using two (contrived) connections. The first connection does not change the fiber coordinate during the parallel transport of an element of the fiber bundle (Fig. 5a). Such a connection is said to be flat. The second connection, in general, changes the fiber coordinate during the parallel transport of an element of the fiber bundle (Fig. 5b). Such a connection is said to be curved. In particular, the connection in Fig. 5b increases the fiber coordinate in proportion to the unwrapped winding angle of a path about point $P$. The change in fiber coordinate after parallel transport of an element of the fiber bundle around a loop in the base manifold is known as holonomy ($\gamma$ in Fig. 5b).
Figure 5: Parallel transport on the fiber bundle of Fig. 3 according to a connection. $C$ and $L$ are a curve and a loop on the base manifold, respectively. (a) Fiber bundle with a flat connection. Parallel transport of an element $u_{1}$ along $C$ does not change the fiber coordinate $f_{1}$. It follows that parallel transport of an element $u_{2}$ along the loop $L$ does not change it. (b) Fiber bundle with a curved connection. Parallel transport of an element $u_{1}$ along $C$ changes the fiber coordinate from $f_{1}$ to $f_{1}^{\prime}$. Parallel transport of an element $u_{2}$ along the loop $L$ changes the fiber coordinate from $f_{2}$ to $f_{2}^{\prime}$, resulting in a holonomy of $\gamma$.
@@
Figure 6a illustrates parallel transport on a bundle constructed by attaching a perpendicular line $\mathbb{R}$ at each point on the plane $\mathbb{R}^{2}$. The bundle is known as the normal bundle over the plane, and it is denoted as $\mathcal{N}\mathbb{R}^{2}$. Its base manifold is $\mathbb{R}^{2}$, and the fiber is $\mathbb{R}$. As the base manifold is a Euclidean space, $\mathcal{N}\mathbb{R}^{2}$ is trivial. $\mathcal{N}\mathbb{R}^{2}$ is, in fact, equal to $\mathbb{R}^{3}$ by construction. A point $u$ of $\mathcal{N}\mathbb{R}^{2}$ describes a vector $v$ normal to the plane $\mathbb{R}^{2}$ that is attached at a point $\mathbf{R}$ of $\mathbb{R}^{2}$. Its fiber coordinate is the vector $v$ and the base coordinate is the point $\mathbf{R}$. Parallel transport on $\mathcal{N}\mathbb{R}^{2}$ corresponds to moving normal vectors along a curve on the plane without rotating them or changing their lengths. The fiber coordinate does not change along the path upon parallel transport (Fig. 6a), indicating that $\mathcal{N}\mathbb{R}^{2}$ is flat. The parallel transport of the fiber elements along any closed loop such as $\mathbf{R}_{1}\mathbf{R}_{2}\mathbf{R}_{3}\mathbf{R}_{4}$ does not result in a holonomy.
@@
The above characterization explains the different mechanisms required to obtain a holonomy with either geometric or topological characteristics. Holonomy with a geometric nature depends on local features of the fiber bundle, namely the curvature induced by the connection, which can be computed pointwise [60]. Holonomy with a topological nature depends on a global feature of the fiber bundle, namely the nontrivial nature of the fiber bundle. Note that trivial and nontrivial fiber bundles are locally indistinguishable because local neighborhoods of any fiber bundle are equivalent to Euclidean spaces. In general, quantities related to the fiber bundle that continuously change with perturbations (geometric nature) depend on local features, and quantities that are invariant to perturbations (topological nature) depend on global features.
@@ 5.1 Effect on the overall phase
Consider a quasi-1D semi-infinite waveguide $W$ whose cross sectional properties are described by a set of $n$ parameters, $\mathbf{R}(z)$ (Fig. 7a). The parameters vary adiabatically in a cycle along the longitudinal direction ($z$ axis). The space of the parameter values is generally equivalent to $\mathbb{R}^{n}$. Each parameter value $\mathbf{R}$ can define a uniform waveguide and its guided modes. Recall from Sec. 3 that a parameter value $\mathbf{R}_{0}$ resulting in a waveguide with fully degenerate modes is said to be a full degeneracy. The parameter space generally consists of nondegenerate values; only specific parameter values $P$ are fully degenerate. The nondegenerate parameter values defines the base manifold $\mathcal{M}$, that is, $\mathcal{M}=\mathbb{R}^{n}-P$ (Figs. 7b-d). A parameter space without full degeneracies results in a Euclidean base manifold and leads to a trivial fiber bundle (Fig. 7b). A parameter space with full degeneracies generally leads to a nontrivial fiber bundle (Fig. 7d).
@@
Figure 7: The geometric phase in an elastic waveguide is modeled using a fiber bundle. (a) A quasi-one-dimensional waveguide $W$ with cross sectional properties $\mathbf{R}$ varying in an adiabatic cycle along $z$. The waveguide is excited by a guided mode eigenfunction $\mathbf{U(x,y,\mathbf{R}(0))}$ at $z=0$. The disturbance propagates to $z=z_{0}$, where the response acquires a geometric phase $\gamma$. (b) Base manifold $\mathcal{M}=\mathbb{R}^{2}$ and fiber $U(1)$ resulting in the fiber bundle $\mathcal{E}_{1}$. Waveguide $W$ corresponds to the loop $L$ on the base manifold, which is parameterized by the $z$ coordinate. The fiber $U(1)$ is represented by unwrapping it into a line, where the color denotes the fiber coordinate. The fiber is attached at each point, as shown at $\mathbf{R}(z_{0})$, to create the fiber bundle $\mathcal{E}_{1}$. The colors on the fiber bundle also correspond to the fiber coordinate. Parallel transport of $u$ along $L$ models wave propagation in $W$. It results in a holonomy $\gamma$, which is a nontopological geometric phase. (c) Base manifold $\mathcal{M}=\mathbb{R}^{2}-\{\mathbf{0}\}$ and fiber $\{0,\pi\}$ resulting in a trivial fiber bundle $\mathcal{E}_{2}$. The fiber $\{0,\pi\}$, depicted by blue cubes, is attached at a point $\mathbf{R}(z_{0})$. (d) Base manifold $\mathcal{M}=\mathbb{R}^{2}-\{\mathbf{0}\}$ and fiber $\{0,\pi\}$ resulting in a nontrivial fiber bundle $\mathcal{E}_{3}$. Two loops $L_{1}$ and $L_{2}$ are marked. Parallel transport of an element $u$ along $L_{1}$ ($L_{2}$) results in a topological geometric phase (no geometric phase). (e) The edge of a Möbius strip, obtained by restricting $\mathcal{E}_{3}$ to loop $L_{1}$. Parallel transport of $u$ is highlighted. (f) The edges of a cylinder, obtained from restricting $\mathcal{E}_{3}$ to loop $L_{2}$. Parallel transport of $u$ is highlighted.
@@
The waveguide $W$ corresponds to a closed loop $L$ on the base manifold, which is parameterized by the longitudinal coordinate $z$ (Fig. 7b). Let the waveguide be harmonically excited with angular frequency $\omega$ triggering a flexural mode whose eigenfunction is $\mathbf{U}(x,y;\mathbf{R}(0))$ at $z=0$. The steady state displacement field at a location $z_{0}$ is determined up to the geometric phase as (Sec. 3)
@@
The geometric phase $\gamma$ accounts for the phase ambiguity of the guided mode eigenfunction (Sec. 2), that is, if $\mathbf{U}(x,y;\mathbf{R}(z_{0}))$ is an eigenfunction, then so is $\mathbf{U}(x,y;\mathbf{R}(z_{0}))e^{\mathrm{i}\gamma}$. In general, the geometric phase $\gamma$ may assume any value from 00 to $2\pi$, that is, any point on the complex unit circle $U(1)$. The geometric phase is described as the fiber coordinate on the fiber $U(1)$ attached to the point $\mathbf{R}$ of the base manifold. For ease of visualization, the fiber is unwrapped into a line whose color indicates the fiber coordinate (inset of Fig. 7b). The base manifold and fibers give rise to a fiber bundle (Fig. 7b). In systems with time reversal symmetry, such as conservative elastic waveguides, the guided modes are described by real-valued eigenfunctions. Thus, a computed eigenfunction is ambiguous only in its sign, and the geometric phase is restricted to $\{0,\pi\}$. The geometric phase is described by the fiber $\{0,\pi\}$. Some possible fiber bundles are shown in Figs. 7c,d. In this manner, the response of a waveguide to a guided mode excitation has been modeled as a fiber bundle with base manifold as the space of nondegenerate parameters $\mathbf{R}$ and the fiber as $U(1)$ or $\{0,\pi\}$.
A connection on this fiber bundle is provided by the governing equations of motion and the adiabatic theorem. The adiabatic theorem limits the allowable magnitude of change in the guided mode eigenfunction over a small longitudinal distance $\Delta z$ to be of the order of $(\Delta z)^{2}$. That is,
@@
Equation (14) defines a connection called the adiabatic connection or the Berry connection [1, 18]. If the fiber is $\{0,\pi\}$, the adiabatic connection simply ensures that the fiber coordinate does not change abruptly (from 00 to $\pi$ or vice versa) as the base point moves in a neighborhood equivalent to the Euclidean space, hence implying that the connection is flat.
@@
Examples of such geometric phase in elastic waves can be found in Refs. [49, 45].
If the fiber bundle is nontrivial with fiber $\{0,\pi\}$, the geometric phase is topological. This situation is exemplified by the discussions on flexural waves propagating in 1D waveguides with adiabatically-varying triangular cross section (Sec. 3.1). The waveguide property varying along the longitudinal direction is a geometric perturbation $\mathbf{R}=(\Delta x,\Delta y)$ to the cross section. The perturbation parameters lead to a two-dimensional parameter space with a full degeneracy at $\mathbf{R}=(0,0)$ arising from the symmetry of the corresponding uniform waveguide, which has an equilateral triangular cross section. Thus, the base manifold is the punctured plane $\mathbb{R}^{2}-\{\mathbf{0}\}$, and the fiber is $\{0,\pi\}$. The base manifold and fiber gives rise to one of the following fiber bundles: (i) a stack of two punctured planes $\mathcal{E}_{2}$ (formally $(\mathbb{R}^{2}-\{\mathbf{0}\})\times\{0,\pi\}$; which is trivial) shown in Fig. 7c, or (ii) a thickened Möbius strip-like bundle $\mathcal{E}_{3}$ (nontrivial) visualized in Fig. 7d (c.f. Sec. 4.1). Note that the self-intersection is an artifact of describing the fiber bundle in 3D space. As the simulations of Sec. 3.1 confirm the appearance of a $\pi$-valued geometric phase, the fiber bundle must be nontrivial (because a trivial fiber bundle with flat connection does not exhibit holonomy).
@@ 5.2 Effect on the polarization angle
For this case, the construction of the fiber bundle and the identification of the connection is discussed using the example of polarized flexural waves propagating in a waveguide (Sec. 3.2).
Consider a waveguide $W$ generated by the closed loop $C_{2}$ on the sphere (Fig. 2g,h). Let the path coordinate along the centerline of the waveguide be $s$. Let the waveguide $W$ be excited by a flexural guided mode, assumed to be horizontally polarized. At any location $s$, the displacement field is determined up to the polarization vector $\hat{\mathbf{c}}$ (Sec. 3), which can be any unit vector in $\mathbb{R}^{2}$. It is modeled as the fiber coordinate on the fiber $\mathbb{R}^{2}$ attached to the point $\hat{\mathbf{t}}$ of the sphere. The resulting fiber bundle is equivalent to the tangent bundle of the sphere $TS^{2}$ (Fig. 3e).
@@ 5.3 General considerations
The previous two sections highlighted a few general considerations for a waveguide $W$ with adiabatically varying parameters $\mathbf{R}$ along the propagation coordinate $z$ (or $s$): (i) the propagation of waves resulting from a harmonic excitation is described as a parallel transport on a fiber bundle, (ii) the geometric phase is the holonomy arising from this parallel transport. The classification of holonomies according to the nature of the fiber bundle (trivial or nontrivial) and curvature of the connection (flat or curved) applies to geometric phases. In particular, a waveguide exhibiting a topological geometric phase has full degeneracies in the parameter space and is described by a nontrivial fiber bundle analogous to the edge of a Möbius strip. A waveguide exhibiting a nontopological geometric phase is described by a fiber bundle with a curved connection. This classification is verified in the examples in Sec. 3.
@@ 6 Topological classification of elastic waveguides
The previous discussion on the geometric phase has illustrated the role of differential geometry and topology in understanding the dynamics of elastic waves. In this section, we show how the relation between geometric phase, topology, and dynamic response can be leveraged to design elastic waveguides.
In particular, we review the role of the geometric phase in the context of the design of elastic topological metamaterials (ETMs). Broadly speaking, elastic metamaterials [64, 65] are typically periodic structures obtained by repeatedly assembling in space a section (called the unit cell) that embeds all the fundamental structural features of the material. The unit cell properties (e.g., geometry, materials, symmetries, etc.) can be tailored to enable a wide range of unconventional static and dynamic responses, such as negative Poisson’s ratio [66], negative refractive index [67], nonreciprocity [68], and cloaking [69, 70], just to name a few. Given the vastness of the design space, it is very challenging to comprehensively explore unit cell designs in search of specific material response. However, topology can be leveraged as an additional tool to support this search. Certain combinations of the unit cell properties, particularly those involving restrictions on spatial and temporal symmetries, lead to a specific class of metamaterials known as ETMs. The waveguiding properties of ETMs are closely connected to the topological properties of the dispersion relations, in a manner that is elucidated in the remainder of this section.
The ability to characterize the material behavior using topological concepts was first identified in the context of band structure analysis of condensed matter systems. The topological classification of abstract spaces and the physical classification of materials were unified in a remarkable chain of discoveries between 1970 and 2000 (see Refs. [24, 26]). These studies showed how naturally occurring crystalline materials could be modeled as fiber bundles, and the corresponding physical properties were linked to topological invariants  [^1] Mathematical parameters that identify global properties of abstract mathematical spaces; in geometry this parameter is the genus. Topological invariants do not change under continuous deformations of the space. For example, a sphere has genus $g=0$. The sphere can be continuously morphed into a parallelepiped, which is also described by $g=0$; therefore the sphere and the parallelepiped are topologically equivalent objects.. As an example, electric polarization was understood as the holonomy of parallel transport [7]. These investigations led to a “periodic table” of topological materials [26]. According to this classification scheme, materials can be categorized based on the topology of their band structure and on the value of the associated invariant. This invariant captures the global (topological) properties of the band structure, hence characterizing the dynamic behavior of the medium at a much deeper level that is independent from specific values of the material parameters. Materials characterized by the same invariant form a topological material class, which means that the topological nature of their band structure is intrinsically equivalent. On the other side, materials in distinct topological classes have inequivalent band structures that are typically separated by degeneracies; more precisely, if a variation of the material parameters changes the topological class, the corresponding band structure cannot be morphed into each other unless the band gap closes and reopens (i.e. degeneracies are created and annihilated). Further, in accordance to the bulk-boundary correspondence principle [24, 26], an interface between two topologically distinct materials, whose invariants differ by $n$ units, supports $n$ localized excitations called edge modes. These edge modes are protected against back-scattering produced by any potential defect or perturbation of the interface because the edge modes are determined by the global (topological) properties of the bulk, not by the local properties of the interface. The only way to alter the edge modes is to destroy the topological property of the bulk. These unique properties of the edge modes have led to technological advances in electronics and information processing [71].
The success of the “periodic table” of (quantum) topological materials spurred new directions such as higher-order topological metamaterials [72], non-Abelian materials [22], and topological defects [73]. Each direction further built upon the connections between the topological characterization of the band structure and the corresponding exotic material behaviors. In each of these new material classes, a material is described as a fiber bundle and relates abstract properties of the fiber bundle to the observable behavior of the material.
@@
In the last decades, researchers have successfully adapted some of these concepts to engineer classical (i.e. non-quantum) topological metamaterials for elastic waves, acoustics, and photonics [39, 42, 35]. In the context of elastic topological metamaterials (ETMs), research focused on the identification and realization of material architectures capable of mimicking the dynamic behavior of their quantum mechanical counterparts. It is important to highlight the foundational difference between quantum and classical materials (whose behavior is ruled by quantum and classical mechanics, respectively) ultimately allows only analogies between the two material systems.
The first generation of ETMs replicated specific classes from the “periodic table” of quantum topological materials [39]. Analogous to their quantum counterparts, such ETMs mapped to fiber bundles and topological invariants, and interfaces between topologically distinct ETMs supported robust, localized edge modes according to the bulk-boundary correspondence. The desirable properties of the edge modes found technological applications. For example, in a 2D ETM, a line interface between topologically distinct ETMs supported localized propagating waves immune to back scattering even in the presence of inhomogeneities and defects [33]. Soon after, ETMs based on higher-order topology [74] and topological defects appeared [75].
@@
The initial development of ETMs was rapid since most designs emulated the design characteristics of their quantum counterparts. Nevertheless, researchers have only begun exploring ETM designs and capabilities [76, 41]. To support the development of ETMs, design tools rooted in differential geometry and topology are critical. However, currently these ideas are spread out over different fields and making their role in the design process of ETMs not always directly accessible. This section presents an attempt to fill this gap by using the notions introduced in Sec. 4. In particular, it highlights the link between waveguides, fiber bundles, topological invariants, and dynamics using simple (yet generalizable) examples.
@@
Consider the adiabatically-varying triangular waveguides presented in Sec. 3.1. A waveguide $W$ can be visualized as a closed loop, say $L_{1}$, on the base manifold $\mathbb{R}^{2}-\{\mathbf{0}\}$ (Fig. 8a), as discussed earlier. $L_{1}$ is parameterized by the longitudinal coordinate $z$. A waveguide can also be visualized in terms of the dynamical properties of its cross section. Each point $\mathbf{R}$ on the loop $L_{1}$ corresponds to a cross section of the waveguide $W$ at some coordinate $z$ along the longitudinal axis. This cross section, with parameters $\mathbf{R}(z)$, defines a uniform waveguide denoted $W_{\mathbf{R}(z)}$. The dynamical properties of the flexural guided modes of the uniform waveguide $W_{\mathbf{R}(z)}$ are described by the dispersion relations $\omega(k,\mathbf{R}(z))$. By allowing $z$ to vary, $\omega(k,\mathbf{R}(z))$ can capture the dispersion relations of every cross section of $W$. For a convenient visualization, we fix $k$ and plot $\omega$ as a function of $z$ in Fig. 8b-d. This plot is referred in the following as the ‘$\omega$-$z$’ plot. The two curves in a $\omega$-$z$ plot generally do not intersect. The curves intersect if, and only if, the flexural guided modes for some $W_{\mathbf{R}(z)}$ are fully degenerate, which means that the loop corresponding to $W$ passes through the full degeneracy (the origin). This situation is illustrated by loop $L_{2}$ in Fig. 8a and the corresponding $\omega$-$z$ plot in Fig. 8c. We digress to note that while the point degeneracy on plots analogous to $\omega$-$z$ curves is crucial in the context of topological classification, its connection with a full degeneracy on the dispersion relation is a peculiarity in the present system. In general, full degeneracies are rare in the general context of topological classification.